OpenStudy (anonymous):

Which of the following is an equivalent form of the following solution? 2x-3y=3 a.) y= -2/3x-1 b.) y= 3/2x-1 c.) y= 2/3x-1 d.) y= 3x-1

OpenStudy (anonymous):

we have 2x - 3y = 3 first shift 2x to right hand side for this subtract 2x from both sides so we get 2x - 2x - 3y = 3 - 2x ==> 0 - 3y = 3 - 2x so we get -3y = 3 - 2x now we need to get rid of -3 from left side so to get rid -3 ( here note that -3 is in multiplication operation with yso we need to divide both side by -3) so divide both side by -3 so we get y = (3/-3) - (2x/3) = (-1 - 2x/3) therfore we get y = -2x/3 - 1

OpenStudy (anonymous):

have u understood ?